Circular No. 6895

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980425 T. J. Galama and P. M. Vreeswijk, University of Amsterdam; E. Pian

and F. Frontera, CNR, Bologna; and V. Doublier and J.-F. Gonzalez, European Southern Observatory (ESO), report: “Comparison of ESO New Technology Telescope images obtained on Apr. 28.4 and May 1.3 UT shows that a point source in the BeppoSAX WFC error box of GRB 980425 (IAUC 6884) not visible in the Digitized Sky Survey brightened by 0.7 mag from R = 15.7 to 15.0 (± 0.1). The object is located at α = 19^{h}35^{m}03^{s}.31, δ = −52^{o}50′44′′.8 (equinox 2000.0), offset from the nucleus of the face-on spiral galaxy (ESO 184-G82). Its position does not coincide with either of the two x-ray sources in the error box of GRB 980425 (Pian et al. 1998, GCN 61), and it is therefore not clear whether the source is related to GRB 980425, or whether it is, e.g., a supernova. A finding chart for it can be found at

http://www.astro.uva.nl/titus/grb980425.html.”

C. Lidman, V. Doublier, J.-F. Gonzalez, T. Augusteijn, O. R. Hainaut, H. Boehnhardt, F. Patat, and B. Leibundgut, ESO, write: “We have observed the supernova candidate discovered by Galama et al. (see above). The object is located in a spiral arm of the barred spiral galaxy ESO 184-G82, the redshift of which was measured at 2528 km/s (heliocentric) from a spectrum obtained with the 3.6-m NTT (+ EMMI). The object displays a steep magnitude increase, as indicated by the following photometry (± 0.05 mag) with the NTT (Danish 1.54-m telescope on May 3): Apr. 28.4 UT, R = 15.77; May 1.3, R = 14.83; 3.30–3.38, V = 14.81, R = 14.35, I = 14.40; 4.4, V = 14.23, R = 14.28, I = 14.27; 6.4, V = 14.00, R = 14.04, I = 14.30. With the ESO/MPI 2.2-m telescope (+ IRAC2) on May 6.4, we obtained the following preliminary photometry J = 11.5, H = 11.6; K = 11.9. A finding chart and secondary photometric standards are avail-

able at http://sc6.sc.eso.org/~ohainaut/SN. Spectra were obtained on

May 1.4 (NTT), 3.4 (1.5-m Danish telescope + DFOSC), 4.4 (ESO 3.6-m telescope + EFOSC2), and 6.4 (NTT). Apart from Hα (probably from the galaxy), the mostly featureless spectra display some broad lines in the range 350–500 nm, then a steep decrease over 500–700 nm, and a plateau from 700 to 1000 nm, with very broad bumps at 620 and 800 nm (spectra of the region 350–900 nm are displayed at the same URL). The relative intensity of the different regions of the spectrum is changing from day to day. The absence of H lines suggests that the object is not a type-II supernova; the lack of Si at 615 nm indicates that it is not a regular type-Ia supernova. The nature of this puzzling object still evades identification, as does its relation to GRB 980425 or to the galaxy. The ESO team continues monitoring.”

1998 May 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green